OpenStudy (anonymous):

Determine the recursive formula for the sequence -7, -13, -19, -25, -31....

OpenStudy (anonymous):

x = x + (-6)

OpenStudy (turingtest):

\[x_{n+1}=x_{n}-6\]the subscripts are important

OpenStudy (anonymous):

Thanks for that correction @TuringTest

OpenStudy (anonymous):

think this squence like a loop :D then by every step it rises i++ x = x - 6, but TuringTest's answer is more good for math :P just old habits from programming lol